OpenStudy (anonymous):

What is the simplified form of 4/3y - 5/y^2

OpenStudy (anonymous):

a) 5y-12/3y^2 b) 4y-14/3y^2 c) 12-5y/3y^2 d) 15-4y/3y^2

OpenStudy (anonymous):

@sarahusher this is just like the other one right?

OpenStudy (anonymous):

well firstly we know that: \[\frac{ a }{ b } - \frac{ c }{ d } = \frac{ a }{ b } + \frac{ d }{ c }\] then: \[\frac{ a }{ b } + \frac{ c }{ d } = \frac{ a*d + c*b }{ b*d }\] so \[\frac{ 4 }{ 3y } - \frac{ 5 }{ y ^{2} } = \frac{ 4 }{ 3y } + \frac{ y ^{2} }{ 5 }\] yes! @danny562 , this is exactly like the other one!

OpenStudy (rishabh.mission):

(4)/(3y)-(5)/(y^(2)) = (4y)/(3y^2) - 5(3) / (3y^2) = (4y-15) / (3y^2)

OpenStudy (anonymous):

@danny562 , have i written the correct fractions above? I get a different answer to those you give

OpenStudy (anonymous):

@satellite73 Can you see what you get please?